# KScope: A Framework for Characterizing the Knowledge Status of Language Models

### v1

@@ Abstract
Characterizing a large language model’s (LLM’s) knowledge of a given question is challenging. As a result, prior work has primarily examined LLM behavior under knowledge conflicts, where the model’s internal parametric memory contradicts information in the external context. However, this does not fully reflect how well the model knows the answer to the question. In this paper, we first introduce a taxonomy of five knowledge statuses based on the consistency and correctness of LLM knowledge modes. We then propose KScope, a hierarchical framework of statistical tests that progressively refines hypotheses about knowledge modes and characterizes LLM knowledge into one of these five statuses. We apply KScope to nine LLMs across four datasets and systematically establish: (1) Supporting context narrows knowledge gaps across models. (2) Context features related to difficulty, relevance, and familiarity drive successful knowledge updates. (3) LLMs exhibit similar feature preferences when partially correct or conflicted, but diverge sharply when consistently wrong. (4) Context summarization constrained by our feature analysis, together with enhanced credibility, further improves update effectiveness and generalizes across LLMs. [^1] Our code is available at https://github.com/xiaoyuxin1002/KScope.
@@ 1 Introduction
However, prior studies on knowledge conflicts do not fully characterize an LLM’s underlying knowledge of a given question. Recent work usually represents LLM’s knowledge via the most likely response (kortukov2024studying, ; wang2024resolving, ; xie2024adaptive, ), which may overlook the coexistence of multiple competing modes in an answer distribution. Moreover, entropy-based uncertainty metrics (du2024context, ; marjanovic2024dynamicqa, ) capture overall uncertainty instead of mode structure, and would assign similar entropy values (approximately $1.37$) to distributions [0.45,[0.45,[ 0.45 , 0.45,0.1]0.45,0.1]0.45 , 0.1 ] and $[0.6,0.2,0.2]$. Yet the first distribution reflects conflicting knowledge, while the second shows consistent preference.
@@
We evaluate the knowledge status of nine LLMs across four datasets (Section 5), and find that supporting context increases the proportion of consistent correct knowledge across all datasets and models. We note that the two healthcare-related datasets (Hemonc (hemonc, ) and PubMedQA (jin2019pubmedqa, )) exhibit lower levels of consistent correct knowledge than the general-domain datasets (NQ (kwiatkowski2019natural, ) and HotpotQA (yang2018hotpotqa, )), both before and after providing context. Larger LLMs within each model family exhibit higher proportions of consistent correct parametric knowledge. Among them, Llama-3 (grattafiori2024llama, ) achieves the highest proportion, followed by Qwen-2.5 (yang2024qwen2, ) and Gemma-2 (team2024gemma, ), although these differences narrow when context is introduced.
@@
We summarize our contributions in this paper as follows:
@@
  • We validate that constrained context summarization, combined with improved credibility, substantially boosts successful knowledge updates across all statuses and generalizes well.
@@ 2 Related Work
Knowledge Update. To incorporate accurate and up-to-date information, retrieval-augmented generation systems (asai2024selfrag, ; gao2023retrieval, ; guu2020realm, ; lewis2020retrieval, ; ram2023context, ; shi2024replug, ) retrieve relevant context from external corpora to support LLMs in knowledge-intensive tasks. The retrieved context can be further augmented to enhance the effectiveness of knowledge updates (tan2024blinded, ; wan2024evidence, ; xu2024earth, ). Other approaches involve directly editing a model’s parametric knowledge (fang2025alphaedit, ; hartvigsen2023aging, ; wang2024wise, ; wang2024editing, ) or applying mechanistic interventions at inference time (jin2024cutting, ; li2023inference, ; zhao2024steering, ). In this work, we adopt a gold-retrieval setting and systematically evaluate various context augmentation strategies tailored to each identified knowledge status.
@@ 3.1 LLM Knowledge Status: Consistency and Correctness
To assess the consistency dimension, we examine three possible structures of $\mathcal{Y}_{p}$: (1) $\mathcal{Y}_{p}=\mathcal{Y}$, (2) $1<|\mathcal{Y}_{p}|<|\mathcal{Y}|$, and (3) $|\mathcal{Y}_{p}|=1$, where ||$\cdot$|| ⋅ | denotes the cardinality of a set. We evaluate correctness by checking whether the ground-truth answer $y^{*}\in\mathcal{Y}_{p}$. Based on these two dimensions, we formulate a taxonomy of five parametric knowledge statuses $P=\text{status}(\mathcal{Y}_{p})$, as illustrated in Figure 1, that characterize the knowledge of $f$ with respect to the question $x$:
@@
Figure 3: Characterization results from applying the KScope framework to nine LLMs across four datasets. Overall, most LLMs exhibit the highest proportion of consistent correct parametric knowledge status, which further increases when supporting context is provided.
@@ 4 Experiment Setup
  • Hemonc (hemonc, ) is a healthcare dataset extracted from a regularly maintained oncology reference database. It consists of $6{,}212$ clinical study instances, each comparing the efficacy of a regimen versus a comparator for a given medical condition, labeled as superior, inferior, or no difference. To reduce positional bias, we permute the order of each regimen–comparator pair. The context for each question consists of abstracts from the associated PubMed articles.
  • PubMedQA (jin2019pubmedqa, ) consists of $1{,}000$ medical research questions, each labeled with a yes, no, or maybe answer. The context comes from the corresponding PubMed abstracts.
  • NQ (kwiatkowski2019natural, ) contains $3{,}596$ Google search queries, retrieving Wikipedia pages as context.
  • HotpotQA (yang2018hotpotqa, ) contains $6{,}119$ multi-hop reasoning questions in the general domain, using sentence-level supporting facts extracted from relevant Wikipedia articles as context.
All supporting contexts here contain evidence for the ground-truth answer. To enable direct comparison of LLMs’ knowledge statuses across datasets, we convert NQ and HotpotQA into three-option classification tasks. Following (xu2024earth, ), we prompt GPT-4o (hurst2024gpt, ) to generate two additional wrong options for each question (details in Appendix A). We note that although not evaluated here, KScope remains applicable to open-ended questions and classification with more options, as discussed in Section 3.
@@
5 Q1: How Does Context Update LLMs’ Knowledge Status?
Using the setup in Section 4, we first characterize the distributions of LLMs’ parametric and contextual knowledge statuses (Section 5.1), then analyze how context induces shifts between them (Section 5.2).
5.1 Distribution of Parametric vs. Contextual Knowledge Status
We present the characterization results from applying KScope to the nine LLMs across the four datasets in Figure 3. When relying solely on parametric knowledge (Figure 3 (a)), LLMs exhibit consistent knowledge—whether correct or wrong—more frequently than conflicting knowledge. When conflicting knowledge occurs, the correct answer is usually included among the knowledge modes. Some outliers show a higher proportion of absent knowledge, such as Gemma-2B on Hemonc.
When LLMs are provided with supporting context (Figure 3 (b)), the proportion of consistent correct knowledge status significantly increases across all datasets and models, with the largest improvement in HotpotQA and the smallest in Hemonc. This highlights the effectiveness of retrieval-augmented generation (asai2024selfrag, ; gao2023retrieval, ; guu2020realm, ; lewis2020retrieval, ), which aims to enhance LLMs’ knowledge with relevant external information. However, in some cases, context may confuse LLMs, leading them to guess randomly and increasing the proportion of absent knowledge. For example, this occurs with Qwen-3B on PubMedQA and the Gemma family on NQ, likely due to longer context lengths in these datasets—an issue we further investigate in Section 6.
5.2 Context-Induced Shifts in Knowledge Status
We further examine how context shifts the distribution of each knowledge status. Figure 4 (a) presents dataset-level shifts, where distributions of knowledge status are averaged across the nine LLMs. We observe that Hemonc and PubMedQA exhibit lower proportions of consistent correct knowledge than NQ and HotpotQA, both before and after context is provided. Additionally, the supporting context slightly increases the ratio of conflicting correct knowledge in Hemonc and absent knowledge in PubMedQA, highlighting the challenges introduced by medical-domain context (busch2025current, ; singhal2023large, ; thirunavukarasu2023large, ; wang2023large, ).
Within each LLM family, we find that larger models consistently exhibit higher proportions of consistent correct knowledge, both before and after providing context (Figure 4 (b)) (allen2025physics33, ; lu2024scaling, ). Llama achieves the highest proportion of consistent correct parametric knowledge, followed by Qwen and Gemma, although this gap narrows once context is introduced. Appendix B details how context shifts each LLM’s knowledge among different statuses on each dataset.
@@ 6 Q2: What Context Features Drive the Desired Knowledge Update?
Based on the evaluation results in Section 5, we seek to understand what context features drive increases in consistent correct knowledge. We introduce three categories of context features in Section 6.1, and inspect feature importance for each knowledge status in Section 6.2.
@@ 7 Q3: What Context Augmentations Work Best Across Knowledge Statuses?
In this section, we leverage the insights from our feature importance analysis in Section 6 to improve knowledge updates in LLMs.
@@ 8 Discussion
Limitations. Our feature importance analysis in Section 6 focuses on eleven features across three categories. However, it remains underexplored how more nuanced stylistic context features (du2024context, ; wan2024evidence, ; xu2024earth, ) impact LLM knowledge status. Although the KScope framework supports open-ended questions (kamalloo2023evaluating, ; wen2024perception, ) and classification tasks with any number of options (tabatabaei2025can, ; zhou2024quest, ), we restrict our experiments to three-option classification due to computational constraints. We also focus on settings where the input context always includes supporting evidence for the ground truth. This allows us to generate practical insights into how such context updates LLM knowledge to the consistent correct status. In real-world retrieval-augmented generation systems (asai2024selfrag, ; gao2023retrieval, ; guu2020realm, ; lewis2020retrieval, ; ram2023context, ; shi2024replug, ), the retrieved context may include noisy or conflicting information (hou2024wikicontradict, ; su2024textttconflictbank, ), which could introduce additional challenges. We leave the investigation of how context affects LLM knowledge status under these more complex conditions to future work.
@@ Appendix A Additional Details on the Experiment Setup
Among the three existing datasets, PubMedQA [26] uses the MIT license, while NQ and HotpotQA are released under the Apache 2.0 license. All three evaluated LLM families (Gemma [55], Llama [15], and Qwen [72]) are distributed under custom commercial licenses. We run all the experiments in this paper on four NVIDIA A100100100100 GPUs.
@@ Appendix B Additional Results on the Context-Induced Shifts in Knowledge Status
In Section 5, we investigate how context updates LLMs’ knowledge status. Figures 13,  14,  15, and 16 provide detailed breakdowns of how context induces shifts from each parametric knowledge status to contextual knowledge status for each LLM on each dataset.
Figure 9: Hyperparameter search on Hemonc using Llama-8B. KScope achieves stable characterization of LLM knowledge status with $M=20$ question paraphrases (left) and $N=100$ model responses per question. We adopt this configuration in all experiments.
@@ Appendix D Additional Results on the Context Augmentation Strategies
Figure 13: Context-induced shifts in knowledge status for each LLM on the Hemonc dataset.
Figure 14: Context-induced shifts in knowledge status for each LLM on the PubMed dataset.
Figure 15: Context-induced shifts in knowledge status for each LLM on the NQ dataset.
Figure 16: Context-induced shifts in knowledge status for each LLM on the HotpotQA dataset.

### v2

@@ Abstract
Characterizing a large language model’s (LLM’s) knowledge of a given question is challenging. As a result, prior work has primarily examined LLM behavior under knowledge conflicts, where the model’s internal parametric memory contradicts information in the external context. However, this does not fully reflect how well the model knows the answer to the question. In this paper, we first introduce a taxonomy of five knowledge statuses based on the consistency and correctness of LLM knowledge modes. We then propose KScope, a hierarchical framework of statistical tests that progressively refines hypotheses about knowledge modes and characterizes LLM knowledge into one of these five statuses. We apply KScope to nine LLMs across four datasets and systematically establish: (1) Supporting context narrows knowledge gaps across models. (2) Context features related to difficulty, relevance, and familiarity drive successful knowledge updates. (3) LLMs exhibit similar feature preferences when partially correct or conflicted, but diverge sharply when consistently wrong. (4) Context summarization constrained by our feature analysis, together with enhanced credibility, further improves update effectiveness and generalizes across LLMs.
@@ 1 Introduction
However, prior studies on knowledge conflicts do not fully characterize an LLM’s underlying knowledge of a given question. Recent work usually represents LLM’s knowledge via the most likely response (kortukov2024studying, ; wang2024resolving, ; xie2024adaptive, ), which may overlook the coexistence of multiple competing modes in an answer distribution. Moreover, entropy-based uncertainty metrics (du2024context, ; marjanovic2024dynamicqa, ) capture overall uncertainty instead of mode structure, and would assign similar entropy values (approximately $1.37$) to distributions $[0.45,$ $0.45,0.1]$ and $[0.6,0.2,0.2]$. Yet the first distribution reflects conflicting knowledge, while the second shows consistent preference.
@@
We evaluate the knowledge status of nine LLMs across four datasets (Section 5), and find that supporting context increases the proportion of consistent correct knowledge across all datasets and models. In the multi-choice setting, we note that the two healthcare-related datasets (Hemonc (hemonc, ) and PubMedQA (jin2019pubmedqa, )) exhibit lower levels of consistent correct knowledge than the general-domain datasets (NQ (kwiatkowski2019natural, ) and HotpotQA (yang2018hotpotqa, )), both before and after providing context. Within each model family, larger LLMs exhibit higher proportions of consistent correct parametric knowledge. Among them, Llama-3 (grattafiori2024llama, ) achieves the highest proportion, followed by Qwen-2.5 (yang2024qwen2, ) and Gemma-2 (team2024gemma, ), although these differences narrow when context is introduced. Finally, we show that noisy retrieval and open-ended question settings substantially reduce the effectiveness of updating LLMs to exhibit consistently correct knowledge.
@@
We summarize our contributions [^1] Our code is available at https://github.com/xiaoyuxin1002/KScope. in this paper as follows:
@@
  • We validate that constrained context summarization, combined with improved credibility, significantly boosts successful knowledge updates across all statuses and generalizes well.
@@ 2 Related Work
Knowledge Update. To incorporate accurate and up-to-date information, retrieval-augmented generation (RAG) systems (asai2024selfrag, ; gao2023retrieval, ; guu2020realm, ; lewis2020retrieval, ; ram2023context, ; shi2024replug, ) retrieve relevant context from external corpora to support LLMs in knowledge-intensive tasks. The retrieved context can be further augmented to enhance the effectiveness of knowledge updates (tan2024blinded, ; wan2024evidence, ; xu2024earth, ). Other approaches directly edit a model’s parametric knowledge (fang2025alphaedit, ; hartvigsen2023aging, ; wang2024wise, ; wang2024editing, ) or apply mechanistic interventions at inference time (jin2024cutting, ; li2023inference, ; zhao2024steering, ). In this work, we examine knowledge updates under both gold and noisy retrieval settings and systematically evaluate various context augmentation strategies tailored to each knowledge status.
@@ 3.1 LLM Knowledge Status: Consistency and Correctness
To assess the consistency dimension, we examine three possible structures of $\mathcal{Y}_{p}$: (1) $\mathcal{Y}_{p}=\mathcal{Y}$, (2) $1<|\mathcal{Y}_{p}|<|\mathcal{Y}|$, and (3) $|\mathcal{Y}_{p}|=1$, where $|\cdot|$ denotes the cardinality of a set. We evaluate correctness by checking whether the ground-truth answer $y^{*}\in\mathcal{Y}_{p}$. Based on these two dimensions, we formulate a taxonomy of five parametric knowledge statuses $P=\text{status}(\mathcal{Y}_{p})$, as illustrated in Figure 1, that characterize the knowledge of $f$ with respect to the question $x$:
@@
Figure 3: Characterization results from applying the KScope framework to nine LLMs across four datasets. Overall, most LLMs exhibit the highest proportion of consistent correct parametric knowledge status, which further increases when gold context is provided.
@@ 4 Experiment Setup
  • Hemonc (hemonc, ) is a healthcare dataset extracted from a regularly maintained oncology reference database. It consists of 6,2126{,}212 clinical study instances, each comparing the efficacy of a regimen versus a comparator for a given medical condition, labeled as superior, inferior, or no difference. To reduce positional bias, we permute the order of each regimen–comparator pair. The context for each question consists of abstracts from the associated PubMed articles.
  • PubMedQA (jin2019pubmedqa, ) consists of 1,0001{,}000 medical research questions, each labeled with a yes, no, or maybe answer. The context comes from the corresponding PubMed abstracts.
  • NQ (kwiatkowski2019natural, ) contains 3,5963{,}596 Google search queries, retrieving Wikipedia pages as context.
  • HotpotQA (yang2018hotpotqa, ) contains 6,1196{,}119 multi-hop reasoning questions in the general domain, using sentence-level supporting facts extracted from relevant Wikipedia articles as context.
Unless otherwise specified, all the supplied contexts here contain ground-truth evidence. When comparing LLMs’ knowledge statuses in the multi-choice setting, we convert NQ and HotpotQA into three-option classification tasks. Following (xu2024earth, ), we prompt GPT-4o (hurst2024gpt, ) to generate two additional wrong options for each question (details in Appendix A). We note that although not evaluated here, KScope remains applicable to classification with more options, as discussed in Section 3.
@@
Figure 5: Characterization results from applying KScope to nine LLMs on HotpotQA across different settings. Compared to (b), where gold context in the multi-choice setting enables more consistent correct knowledge, (c) noisy context and (e) the open-ended setting yield lower update success. Without context, the Gemma family shows more absent knowledge in (d) the open-ended setting than in (a) the multi-choice setting, whereas the Llama and Qwen families mostly show the opposite trend.
5 Q1: How Does Context Update LLMs’ Knowledge Status?
Using the setup in Section 4, we first characterize the distributions of LLMs’ parametric and contextual knowledge statuses in the multi-choice setting and examine how gold context induces shifts between them in Section 5.1. We further investigate the effects of noisy context in Section 5.2 and analyze knowledge status distributions in the open-ended setting in Section 5.3.
5.1 Knowledge Status in the Multi-Choice Setting with Gold Context
We apply KScope to the nine LLMs across the four datasets in the multi-choice setting and present the results in Figure 3. When relying solely on parametric knowledge (Figure 3 (a)), LLMs exhibit consistent knowledge—whether correct or wrong—more frequently than conflicting knowledge. When conflicting knowledge occurs, the correct answer is usually among the knowledge modes. Some outliers show a higher proportion of absent knowledge, such as Gemma-2B on Hemonc.
When LLMs are provided with gold context (Figure 3 (b)), the proportion of consistent correct knowledge status significantly increases across all datasets and models, with the largest improvement in HotpotQA and the smallest in Hemonc. This highlights the effectiveness of retrieval-augmented generation (asai2024selfrag, ; gao2023retrieval, ; guu2020realm, ; lewis2020retrieval, ), which aims to enhance LLMs’ knowledge with relevant external information. However, in some cases, context may confuse LLMs, leading them to guess randomly and increasing the proportion of absent knowledge. For example, this occurs with Qwen-3B on PubMedQA and the Gemma family on NQ, likely due to longer context lengths in these datasets—an issue we further investigate in Section 6.
We further examine how gold context shifts the distribution of each knowledge status. Figure 4 (a) presents dataset-level shifts, where distributions of knowledge status are averaged across the nine LLMs. We observe that Hemonc and PubMedQA exhibit lower proportions of consistent correct knowledge than NQ and HotpotQA, both before and after context is provided. Additionally, the gold context slightly increases the ratio of conflicting correct knowledge in Hemonc and absent knowledge in PubMedQA, highlighting the challenges introduced by medical-domain context (busch2025current, ; singhal2023large, ; thirunavukarasu2023large, ; wang2023large, ).
Within each LLM family, we find that larger models consistently exhibit higher proportions of consistent correct knowledge, both before and after providing context (Figure 4 (b)) (allen2025physics33, ; lu2024scaling, ). Llama achieves the highest proportion of consistent correct parametric knowledge, followed by Qwen and Gemma, although this gap narrows once context is introduced. Appendix B details how gold context shifts each LLM’s knowledge among different statuses on each dataset in the multi-choice setting.
@@ 5.1 Knowledge Status in the Multi-Choice Setting with Gold Context
5.2 Knowledge Status in the Multi-Choice Setting with Noisy Context
To investigate LLMs’ knowledge status under more realistic retrieval conditions, we apply KScope to the nine LLMs using the top ten Wikipedia paragraphs retrieved for each HotpotQA question, which may or may not include the gold supporting context. As shown in Figure 5 (c), noisy context results in a much lower success rate of updating models to consistent correct knowledge compared to gold context in (b), highlighting the importance of retrieval quality in RAG systems (asai2024selfrag, ; gao2023retrieval, ; guu2020realm, ; lewis2020retrieval, ; ram2023context, ; shi2024replug, ). When the retrieved noisy context lacks evidence for the ground-truth answer, models either refuse to answer, leading to more absent knowledge, or are misled into producing consistently incorrect answers. More details on noisy context-induced shifts in knowledge status are in Appendix B.
5.3 Knowledge Status in the Open-Ended Setting with Gold Context
To demonstrate the generalizability of KScope to open-ended questions (kamalloo2023evaluating, ; wen2024perception, ), we apply it to characterize the knowledge status of the nine LLMs on HotpotQA, this time without providing any pre-defined answer options. Specifically, we generate responses per question and semantically cluster kuhn2023semantic them using gemma-2-9b-it (team2024gemma, ). Based on the size of clusters and the number of invalid answers, we follow the procedure in Section 3.3 to infer the LLM’s knowledge status.
As shown in Figure 5 (d), knowledge status distributions vary notably across model families. Without pre-defined options or contextual support, the Gemma family often refuses to answer, leading to a higher proportion of absent knowledge compared to the multi-choice setting in (a). In contrast, Llama and Qwen show a substantial increase in the consistent correct knowledge under these conditions, with the exception of Qwen-14B. Gold context still significantly boosts consistent correct knowledge in the open-ended setting in (e), though the improvement is smaller than in the multi-choice setting in (b). Appendix B provides more details on the shift in knowledge status in the open-ended setting.
@@ 6 Q2: What Context Features Drive the Desired Knowledge Update?
Based on the results in Section 5, we seek to understand what context features drive increases in consistent correct knowledge. We introduce three categories of context features in Section 6.1, and inspect feature importance for each knowledge status in Section 6.2. To enable direct comparison across datasets, we focus on the multi-choice setting with gold context in the following sections.
@@ 7 Q3: What Context Augmentations Work Best Across Knowledge Statuses?
In this section, we leverage the insights from our feature importance analysis in Section 6 to improve knowledge updates in LLMs. We again focus on the multi-choice setting with gold context.
Figure 8: Absolute change ($\%$) in the success rate of knowledge updates for each context augmentation strategy, relative to the original context and averaged across datasets. Integrating credibility metadata into constrained summarization improves the success rate by $4.3\%$ on average across LLMs and statuses, and generalizes well to GPT-4o.
@@ 8 Discussion
Limitations. Our feature importance analysis in Section 6 focuses on eleven features across three categories. However, it remains underexplored how more nuanced stylistic context features (du2024context, ; wan2024evidence, ; xu2024earth, ) impact LLM knowledge status. Although the KScope framework supports classification tasks with any number of options (tabatabaei2025can, ; zhou2024quest, ), we restrict our experiments to three-option classification due to computational constraints. We also experiment with real-world noisy retrieval, but such context can include conflicting information (hou2024wikicontradict, ; su2024textttconflictbank, ) in practical RAG systems (asai2024selfrag, ; gao2023retrieval, ; guu2020realm, ; lewis2020retrieval, ; ram2023context, ; shi2024replug, ), posing additional challenges. We leave the investigation of how context affects LLM knowledge status under these more complex conditions to future work.
@@ Appendix A Additional Details on the Experiment Setup
Among the three existing datasets, PubMedQA [26] uses the MIT license, while NQ and HotpotQA are released under the Apache 2.0 license. All three evaluated LLM families (Gemma [55], Llama [15], and Qwen [72]) are distributed under custom commercial licenses. We run all the experiments in this paper on four NVIDIA A100100 GPUs.
@@ Appendix B Additional Results on the Context-Induced Shifts in Knowledge Status
In Section 5, we investigate how context updates LLMs’ knowledge status. Figures 14,  15,  16, and 17 provide detailed breakdowns of how gold context induces shifts from each parametric knowledge status to contextual knowledge status for each LLM on each dataset in the multi-choice setting. Figure 18 shows the shifts induced by noisy context on HotpotQA in the multi-choice setting, while Figure 19 illustrates the shifts induced by gold context on HotpotQA in the open-ended setting.
Figure 10: Hyperparameter search on Hemonc using Llama-8B with gold context in the multi-choice setting. KScope achieves stable characterization of LLM knowledge status with $M=20$ question paraphrases (left) and $N=100$ model responses per question (right). We adopt this configuration in all experiments.
@@ Appendix D Additional Results on the Context Augmentation Strategies
Figure 14: Shifts in knowledge status induced by gold context for each LLM on the Hemonc dataset in the multi-choice setting.
Figure 15: Shifts in knowledge status induced by gold context for each LLM on the PubMed dataset in the multi-choice setting.
Figure 16: Shifts in knowledge status induced by gold context for each LLM on the NQ dataset in the multi-choice setting.
Figure 17: Shifts in knowledge status induced by gold context for each LLM on the HotpotQA dataset in the multi-choice setting.
Figure 18: Shifts in knowledge status induced by noisy context for each LLM on the HotpotQA dataset in the multi-choice setting.
Figure 19: Shifts in knowledge status induced by gold context for each LLM on the HotpotQA dataset in the open-ended setting.